OpenStudy (anonymous):

i just need make sure im right Multiply. 6(4x + 4) A. 48x B. 24x + 24 C. 10x + 10 D. 10x + 24

OpenStudy (anonymous):

B

OpenStudy (anonymous):

hum so i was right..

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

much obliged

OpenStudy (whpalmer4):

This is just the distributive property in action. \[6(4x+4) = 6*4x + 6*4 = 24x + 24\] \[a(b+c) = a*b + a*c\]Also, \[a(b-c) = a*b - a*c\]

OpenStudy (anonymous):

mk cool @whpalmer4